# ImplicitQA: Going beyond frames towards Implicit Video Reasoning

### v1

@@ 4 Benchmarking
Model Lateral Spatial Reasoning Vertical Spatial Reasoning Relative Depth and Proximity Viewpoint and Visibility Motion & Traj. Dynamics Causal and Motivational Reasoning Inferred Counting Physical & Env. Context Social Interaction & Relations Avg. Macro Avg.
Human Baseline 85.4 79.1 80.4 90.0 91.9 94.4 53.9 83.3 100.0 83.0 84.3
@@
LLaVA-Next-Video[20] 36.0 29.6 30.1 48.8 36.3 39.0 30.2 35.7 51.7 33.9 37.5
LLaVA-OneVision[5] 37.3 46.8 35.0 56.1 57.1 57.3 23.3 50.0 55.2 43.4 46.4
LLaVA-Video[21] 36.0 44.0 31.6 56.1 60.4 62.2 14.0 50.0 62.1 42.1 46.3
Qwen2-VL[14] 39.8 46.8 40.6 51.2 52.7 58.5 16.3 35.7 72.4 44.9 46.0
Qwen2.5-VL[1] 41.6 47.2 32.7 61.0 50.5 51.2 25.6 42.9 62.1 42.8 46.1
@@
Our evaluation of 26 VideoQA model configurations across different scales and temporal contexts on the ImplicitQA benchmark reveals several critical insights. Detailed table is presented in Supplementary Section 2. We evaluate a broad range of open-source and proprietary multimodal models on the ImplicitQA benchmark, focusing on their ability to perform implicit reasoning over videos. The evaluation includes multiple model families, scales, and video context lengths. We include the following models in our evaluation: (1) Open-source models: LLaVA-NeXT-Video[20], LLaVA-OneVision[5], LLaVA-Video[21], Qwen2 VL[14], Qwen2.5 VL[1] (2) Closed-source models: GPT-4.1 [10]-full, mini, and nano variants, and the reasoning based GPT-O3 [11] model. We analyze the performance of these models via category-wise accuracy, and overall average accuracy. We also report the macro-average accuracy, which is the average of category-wise accuracies giving equal weight to each category. Additionally, we explore the impact of integrating an explicit reasoning prompt with GPT models. Please refer to the supplementary section 2 for detailed scores at model scale and number of frames.
@@
Human Baseline A non-expert annotator benchmarked themselves against the ImplicitQA dataset using our Implicit-Eval tool. The annotator took around 1 minute per question on average, and achieved an overall score of 83%, with macro average of 84.3%. Inferred Counting proved to be the hardest category, testing the patience of the human annotator and resulting in a score just above 50%. The best models lag behind human baseline the most in lateral spatial reasoning, relative depth and motion dynamics.
@@
Overall Performance Table 2 highlights that the reasoning based GPT O3[11] outperformed all other models significantly, achieving an overall accuracy of 64.1% and a macro average of 68.6%. GPT 4.1[10] also exhibited robust performance, reaching overall accuracy of 54.3%. Open-source models generally performed worse than proprietary models. Early-generation of VideoLLMs like LLaVA-NeXT-Video[20] perform especially poorly, achieving an overall score of 33.9%, indicating limited ability to reason beyond surface-level cues. Current generation open source VideoLLMs perform similarly on the Macro-Average metric (all within a narrow range of 46.0 to 46.4), however they display variation across categories; Qwen models excel at spatial reasoning and viewpoint questions, while LLaVA-OneVision based models (LLaVA-Video is finetuned from LLaVA-OneVision) succeed at Motion, Causal and Contextual Reasoning.
Implicit reasoning remains an unsolved challenge even for leading VideoQA models. More frames or larger models alone are insufficient—we need architectural innovations or novel training paradigms.
@@
Despite being state-of-the-art in traditional VideoQA, no open-source model crosses 50% on ImplicitQA, reflecting the dataset’s challenge and the limitations of current architectures in handling implicit, unstated, or cross-frame reasoning.
@@
Model Capacity vs. Implicit Reasoning Proprietary GPT models notably benefitted from larger scales, with GPT 4.1[10] demonstrating substantial accuracy improvements, indicating that more parameters facilitate effective implicit reasoning. Among open-source models, larger scales like Qwen2.5-VL[1] (32b) consistently outperformed smaller variants (7b and 3b) by a small margin.
@@
(a) Vertical Spatial Reasoning: Among open-weight models, LLaVA-OV [5] performed best with 50.46% accuracy. In proprietary, GPT O3 excelled with a superior accuracy of 72.2%.
(b) Relative Depth and Proximity: GPT-based models demonstrated notable dominance in this category , particularly GPT O3 (55.3%), followed by GPT 4.1[10] (51.1%). Among open-weight models, Qwen2-VL achieved a commendable 40.6%, but significantly lagged proprietary models.
(c) Causal and Motivational Reasoning: This was particularly challenging for all models except GPT variants. GPT O3 notably outperformed all others, achieving 85.4%, significantly ahead of the next best GPT 4.1 at 82.9%. Open-weight models achieved considerably lower performance.
@@
Impact of Reasoning Prompt on GPT Models We introduced a structured reasoning prompt specifically emphasizing spatial relationships and narrative summarization before answering questions. GPT 4.1-Mini[10], with the added reasoning prompt, showed enhanced accuracy across multiple categories compared to the standard GPT 4.1-Mini. Specifically, improvements were evident in vertical spatial reasoning (from 37.96% to 48.61%), viewpoint and visibility (from 39.02% to 56.1%), and motion and trajectory dynamics (from 41.76% to 54.95%) as shown in Figure 8. This clearly indicates that structured prompting focused on spatial and narrative reasoning significantly enhances model performance. Please refer to Supplementary Section 4 for detailed reasoning prompt.
@@
Effect of Number of Frames: Plateauing gains Going from 8 to 16 frames improves performance modestly. Beyond 16 frames to 32 frames, accuracy saturates or slightly improves, but not significantly. For Qwen2.5 VL[1] model it improves 2%.
Uneven performance across reasoning categories Models perform relatively better on Social Interaction and Causal/Motivational Reasoning, possibly reflecting pretraining biases towards human-centric scenarios. Categories involving numerical inference (Inferred Counting) and nuanced spatial reasoning (Relative Depth and Proximity) consistently show lower accuracy, highlighting a need for improvement.
@@ 5 Conclusion
We introduced ImplicitQA, a novel benchmark specifically designed to evaluate implicit reasoning capabilities in VideoQA models. Current benchmarks and methods predominantly address explicit visual understanding, leaving significant gaps in reasoning over narrative-driven, cinematic content, which frequently requires inferring unstated causality, motivations, and relationships. Through extensive experiments involving various models, we demonstrated that existing state-of-the-art exhibits substantial limitations in handling implicit reasoning tasks. Although scaling model sizes and increasing temporal context provide modest improvements, these alone are insufficient to significantly boost performance. Models show particular weaknesses in reasoning categories requiring nuanced numerical inference, detailed spatial relationships, and integration of long-term narrative context. ImplicitQA serves as a rigorous benchmark and a foundational step towards fostering research efforts aimed at bridging the gap between surface-level visual recognition and deeper narrative comprehension, thereby advancing VideoQA towards more human-like understanding and reasoning.
Supplementary Material: ImplicitQA: Going beyond frames towards Implicit Video Reasoning
We organize the supplementary material as follows:
@@ Appendix A Data, Tool and Licenses
We publicly release the benchmark, annotation tool, evaluation scripts and evaluation tool with Apache 2.0 license. The employed multi-modal LMMs are released under Apache 2.0 license. We run all our evaluations on NVIDIA A6000 48GB GPUs, and released the eval scripts for reproducibility.
@@ Appendix B Detailed Results
We present detailed results for varying model scales, temporal context across different models in Table 3. The best-performing open-source models were variants of the Qwen2.5 VL[14] family at 32B scale. Model scale had a noticeable but diminishing effect: Within Qwen2.5 VL, increasing model size from 3B to 32B parameters provided modest performance gains (approximately +2% accuracy improvement). Smaller models, such as LLaVA-OneVision, struggled significantly in challenging reasoning categories irrespective of scale. Distinct performance variations across reasoning categories emerged. Social Interaction was relatively easier, with accuracy up to 79.31% (qwen2.5 VL).
Analysis on increasing frame count generally improves model performance for certain architectures, but this improvement plateaus or slightly degrades beyond a threshold. For instance, LLaVA-NeXT-Video[20] exhibited a peak performance at 16 frames (33.9%) with a marginal decrease at 32 frames (32.56%). Similarly, LLaVA-OneVision[5] performed slightly better at 32 frames (43.16%) compared to 16 frames (43.4%), indicating negligible performance gains. For Qwen2.5-VL[1] 32B model, increasing frames from 4 to 32 resulted in accuracy improvements, particularly evident in the inferred counting & social interaction categories, achieving substantial accuracy improvements from 27.91% to 41.86% and from 62.07% to 79.31%, respectively. This suggests deeper frame context substantially aids in specific implicit reasoning tasks.
@@
Model Scale #Frames Lateral Spatial Reasoning Vertical Spatial Reasoning Relative Depth and Proximity Viewpoint and Visibility Motion & Traj. Dynamics Causal and Motivational Reasoning Inferred Counting Physical & Env. Context Social Interaction & Relations Avg. Macro Avg.
LLaVA-NeXT-Video [20] 7b 8 34.78 31.48 28.95 43.90 37.36 40.24 23.26 42.86 55.17 33.72 37.56
@@ Appendix D Impact of Reasoning Prompt
Model Scale Lateral Spatial Reasoning Vertical Spatial Reasoning Relative Depth and Proximity Viewpoint and Visibility Motion & Traj. Dynamics Causal and Motivational Reasoning Inferred Counting Physical & Env. Context Social Interaction & Relations Avg. Macro Avg.
GPT 4.1 [10] Full 42.90 53.20 51.10 48.80 59.30 82.90 41.90 71.40 75.90 54.30 58.60
GPT 4.1-Reasoning [10] Full 50.90 63.00 51.10 46.30 63.70 79.30 41.90 71.40 86.20 58.20 61.50
GPT 4.1 [10] Mini 44.10 37.96 29.32 39.02 41.76 60.98 32.56 78.57 68.97 40.30 48.14
GPT 4.1-Reasoning [10] Mini 36.02 48.61 33.08 56.10 54.95 65.85 39.53 64.29 72.41 45.07 52.32
@@ Appendix E Experiment Statistical Significance
Lateral Spatial Reasoning Vertical Spatial Reasoning Relative Depth and Proximity Viewpoint and Visibility Motion & Traj. Dynamics Causal and Motivational Reasoning Inferred Counting Physical & Env. Context Social Interaction & Relations Avg.
@@ Appendix F Human Baseline
To establish a reference point for model performance, we evaluated a non-expert human annotator on the ImplicitQA benchmark using our custom-built Implicit-Eval tool. The annotator was not provided with prior exposure to the dataset or answer keys and completed the evaluation at a natural pace, averaging approximately one minute per question. The annotator achieved an overall accuracy of 83.0%, with a macro-average score of 84.3% across reasoning categories. While performance was consistently strong across most categories, Inferred Counting emerged as the most challenging for the human, yielding an accuracy just above 50%. This aligns with the category’s reliance on temporal cues and implicit aggregation of visual information across multiple scenes—factors that often test not only reasoning but also memory and attention.
We further analyze the human accuracy across questions grouped by difficulty level using our hardness-based scoring method. As shown in Figure 17, the non-expert human achieved near-ceiling performance on Easy questions and maintained strong performance on Medium ones. Remarkably, even on Hard questions—defined by consistent failure across models—the human annotator performed well above chance, achieving over 70% accuracy. This analysis underscores the significant gap between human and model capabilities on complex reasoning tasks. Even non-expert humans demonstrate strong generalization and implicit understanding, particularly in scenes that demand spatial, temporal, or causal reasoning—highlighting the limitations of current state-of-the-art Video LMMs.
@@ Appendix H Qualitative Results
To illustrate the complexity and diversity of implicit reasoning required in our benchmark, we present qualitative examples spanning all nine reasoning categories in Figure 21,22,23. For each example, we show the relevant video frames, question, answer choices, correct answer, and model predictions. These examples highlight a wide range of reasoning challenges -from spatial positioning and motion inference to social understanding and causal explanation. Notably, GPT-O3 demonstrates superior performance in most cases.

### v2

@@ 4 Benchmarking
Model Lateral Spatial Reasoning Vertical Spatial Reasoning Relative Depth and Proximity Viewpoint and Visibility Motion & Traj. Dynamics Causal & Motivational Reasoning Inferred Counting Physical & Env. Context Social Interaction & Relations Avg. Macro Avg.
Human Baseline 85.4 79.1 80.4 90.0 91.9 94.4 65.9 83.3 100.0 83.0 85.6
@@
LLaVA-Next-Video 36.0 29.6 30.1 48.8 36.3 39.0 30.2 35.7 51.7 33.9 37.5
LLaVA-OneVision 37.3 46.8 35.0 56.1 57.1 57.3 23.3 50.0 55.2 43.4 46.4
LLaVA-Video 36.0 44.0 31.6 56.1 60.4 62.2 14.0 50.0 62.1 42.1 46.3
Gemma 3 48.5 38.9 32.3 68.3 39.6 59.8 25.6 57.1 58.6 42.1 47.6
InternVL 3 34.8 39.4 37.2 56.1 51.7 64.6 34.9 57.1 75.9 43.3 50.2
Qwen2-VL 39.8 46.8 40.6 51.2 52.7 58.5 16.3 35.7 72.4 44.9 46.0
Qwen2.5-VL 41.6 47.2 32.7 61.0 50.5 51.2 25.6 42.9 62.1 42.8 46.1
@@
Our evaluation of 30 VideoQA model configurations across different scales and temporal contexts on the ImplicitQA benchmark reveals several critical insights. Detailed table is presented in Section B. We evaluate a broad range of open-source and proprietary multimodal models on the ImplicitQA benchmark, focusing on their ability to perform implicit reasoning over videos. The evaluation includes multiple model families, scales, and video context lengths. We include the following models in our evaluation: (1) Open-source models: LLaVA-NeXT-Video Zhang et al. [2024a], LLaVA-OneVision Li et al. [2024a], LLaVA-Video Zhang et al. [2024b], Qwen2 VL Wang et al. [2024b], Qwen2.5 VL Bai et al. [2025], InternVL3 Zhu et al. [2025], Gemma 3 Team et al. [2025], (2) Closed-source models: GPT-4.1 OpenAI [2024b]-full, mini, and nano variants, and the reasoning based GPT-O3 OpenAI [2024a] model, Gemini 2.5 Flash Comanici et al. [2025], Claude 4 sonnet Anthropic . We analyze the performance of these models via category-wise accuracy, and overall average accuracy. We also report the macro-average accuracy, which is the average of category-wise accuracies giving equal weight to each category. Additionally, we explore the impact of integrating an explicit reasoning prompt with GPT models. Please refer to the Section B for detailed scores at model scale and number of frames.
@@
Human Baseline Non-expert annotators benchmarked themselves against the ImplicitQA dataset using our Implicit-Eval tool. The annotators took around 1 minute per question on average, and achieved an overall score of 83%, with macro average of 85.6%. The best models lag behind human baseline the most in lateral spatial reasoning, relative depth and motion dynamics.
@@
Overall Performance Table 3 highlights that the reasoning based GPT O3 outperformed all other models significantly, achieving an overall accuracy of 64.1% and a macro average of 68.6%. GPT 4.1 also exhibited robust performance, reaching overall accuracy of 54.3% followed by Gemini 2.5 Flash and Claude 4 sonnet. Open-source models generally performed worse than proprietary models. Early-generation of VideoLLMs like LLaVA-NeXT-Video perform especially poorly, achieving an overall score of 33.9%, indicating limited ability to reason beyond surface-level cues. Current generation open source VideoLLMs perform similarly on the Macro-Average metric (all within a range of 46.0 to 50.2), however they display variation across categories; Qwen, Gemma, Internvl models excel at spatial reasoning and viewpoint questions, while LLaVA-OneVision based models (LLaVA-Video is finetuned from LLaVA-OneVision) succeed at Motion, Motivational Reasoning.
@@
Implicit reasoning remains an unsolved challenge even for leading VideoQA models. More frames or larger models alone are insufficient, we need architectural innovations or novel training paradigms. Despite being state-of-the-art in traditional VideoQA, no open-source model crosses 50% on ImplicitQA, reflecting the dataset’s challenge and the limitations of current architectures in handling implicit, unstated, or cross-frame reasoning.
@@
Model Capacity vs. Implicit Reasoning GPT models notably benefitted from larger scales, with GPT 4.1 OpenAI [2024b] demonstrating substantial accuracy improvements, indicating more parameters facilitate effective implicit reasoning. Among open-source models, larger scales like Qwen2.5-VL Bai et al. [2025]-32b consistently outperformed smaller variants (7b, 3b) by a small margin.
@@
(a) Vertical Spatial Reasoning: Among open-weight models, LLaVA-OV Li et al. [2024a] performed best with 50.46% accuracy. In proprietary, GPT O3 excelled with a superior accuracy of 72.2%.
(b) Relative Depth and Proximity: GPT-based models demonstrated notable dominance in this category , particularly GPT O3 (55.3%), followed by GPT 4.1 OpenAI [2024b] (51.1%). Among open-weight models, Qwen2-VL achieved a commendable 40.6%, but lagged proprietary models.
(c) Motivational Reasoning: This was particularly challenging for all models except GPT variants. GPT O3 notably outperformed all others, achieving 85.4%, significantly ahead of the next best GPT 4.1 at 82.9%. Open-weight models achieved considerably lower performance.
@@
Impact of Reasoning Prompt on GPT Models We introduced a structured reasoning prompt specifically emphasizing spatial relationships and narrative summarization before answering questions. GPT 4.1-Mini OpenAI [2024b], with the added reasoning prompt, showed enhanced accuracy across multiple categories compared to the standard GPT 4.1-Mini. Specifically, improvements were evident in vertical spatial reasoning (from 37.96% to 48.61%), viewpoint and visibility (from 39.02% to 56.1%), and motion and trajectory dynamics (from 41.76% to 54.95%) as shown in Figure 8. This clearly indicates that structured prompting focused on spatial and narrative reasoning significantly enhances model performance. Please refer to Section D for detailed reasoning prompt.
@@
Effect of Number of Frames: Going from 8 to 16 frames improves performance modestly. From 16 frames to 32 frames, accuracy saturates or slightly improves, but not significantly. For Qwen2.5 VL model it improves 2%.
Uneven performance across reasoning categories Models perform relatively better on Social Interaction and Motivational Reasoning, possibly reflecting pretraining biases towards human-centric scenarios. Categories involving numerical inference (Inferred Counting) and nuanced spatial reasoning (Relative Depth and Proximity) consistently show lower accuracy, highlighting a need for improvement.
@@ 5 Conclusion
We introduce ImplicitQA, a novel benchmark for evaluating implicit reasoning in videos. Prior datasets largely target explicit visual understanding, leaving a gap in reasoning over narrative, cinematic content where answers often depend on inferred motivations and relationships. Across extensive experiments with contemporary models, we find that state-of-the-art systems struggle on implicit reasoning: scaling model size and extending temporal context offer only modest gains. Weaknesses are especially pronounced for categories requiring nuanced numerical inference, fine-grained spatial reasoning, and integration of long-term narrative context. ImplicitQA provides a rigorous evaluation protocol and baseline analyses to catalyze research that moves VideoQA beyond surface-level recognition toward deeper, human-like narrative understanding.
Appendix: ImplicitQA: Going beyond frames towards Implicit Video Reasoning
We organize the appendix as follows:
@@ Appendix A Data, Tool and Licenses
We plan to publicly release the benchmark, annotation tool, evaluation scripts and evaluation tool with Apache 2.0 license upon acceptance. The employed multi-modal LMMs are released under Apache 2.0 license. We ran all our evaluations on NVIDIA A6000 48GB GPUs, and will release the eval scripts for reproducibility.
@@ Appendix B Detailed Results
We present detailed results for varying model scales, temporal context across different models in Table 4. The best-performing open-source models were variants of the Qwen2.5 VLWang et al. [2024b] family at 32B scale. Model scale had a noticeable but diminishing effect: Within Qwen2.5 VL, increasing model size from 3B to 32B parameters provided modest performance gains (approximately +2% accuracy improvement). Smaller models, such as LLaVA-OneVision, struggled significantly in challenging reasoning categories irrespective of scale. Distinct performance variations across reasoning categories emerged. Social Interaction was relatively easier, with accuracy up to 79.31% (qwen2.5 VL).
Analysis on increasing frame count generally improves model performance for certain architectures, but this improvement plateaus or slightly degrades beyond a threshold. For instance, LLaVA-NeXT-VideoZhang et al. [2024a] exhibited a peak performance at 16 frames (33.9%) with a marginal decrease at 32 frames (32.56%). Similarly, LLaVA-OneVisionLi et al. [2024a] performed slightly better at 32 frames (43.16%) compared to 16 frames (43.4%), indicating negligible performance gains. For Qwen2.5-VLBai et al. [2025] 32B model, increasing frames from 4 to 32 resulted in accuracy improvements, particularly evident in the inferred counting & social interaction categories, achieving substantial accuracy improvements from 27.91% to 41.86% and from 62.07% to 79.31%, respectively. This suggests deeper frame context substantially aids in specific implicit reasoning tasks.
@@
Model Scale #Frames Lateral Spatial Reasoning Vertical Spatial Reasoning Relative Depth and Proximity Viewpoint and Visibility Motion & Traj. Dynamics Motivational Reasoning Inferred Counting Physical & Env. Context Social Interaction & Relations Avg. Macro Avg.
LLaVA-NeXT-Video Zhang et al. [2024a] 7b 8 34.78 31.48 28.95 43.90 37.36 40.24 23.26 42.86 55.17 33.72 37.56
@@ C.4 Difficulty
C.5 Question Word Distribution
In Figure 18 we present the word count distribution for questions in ImplicitQA.
Figure 18: ImplicitQA Question word counts
@@ Appendix D Impact of Reasoning Prompt
Model Scale Lateral Spatial Reasoning Vertical Spatial Reasoning Relative Depth and Proximity Viewpoint and Visibility Motion & Traj. Dynamics Motivational Reasoning Inferred Counting Physical & Env. Context Social Interaction & Relations Avg. Macro Avg.
GPT 4.1 OpenAI [2024b] Full 42.90 53.20 51.10 48.80 59.30 82.90 41.90 71.40 75.90 54.30 58.60
GPT 4.1-Reasoning OpenAI [2024b] Full 50.90 63.00 51.10 46.30 63.70 79.30 41.90 71.40 86.20 58.20 61.50
GPT 4.1 OpenAI [2024b] Mini 44.10 37.96 29.32 39.02 41.76 60.98 32.56 78.57 68.97 40.30 48.14
GPT 4.1-Reasoning OpenAI [2024b] Mini 36.02 48.61 33.08 56.10 54.95 65.85 39.53 64.29 72.41 45.07 52.32
@@ Appendix E Experiment Statistical Significance
Lateral Spatial Reasoning Vertical Spatial Reasoning Relative Depth and Proximity Viewpoint and Visibility Motion & Traj. Dynamics Motivational Reasoning Inferred Counting Physical & Env. Context Social Interaction & Relations Avg.
@@ Appendix F Human Baseline
To establish a reference point for model performance, we evaluated with non-expert human annotator on the ImplicitQA benchmark using our custom-built Implicit-Eval tool. The annotator was not provided with prior exposure to the dataset or answer keys and completed the evaluation at a natural pace, averaging approximately one minute per question. The annotator achieved an overall accuracy of 83.0%, with a macro-average score of 85.6% across reasoning categories. While performance was consistently strong across most categories, Inferred Counting emerged as the most challenging for the human. This aligns with the category’s reliance on temporal cues and implicit aggregation of visual information across multiple scenes, factors that often test not only reasoning but also memory and attention.
We further analyze the human accuracy across questions grouped by difficulty level using our hardness-based scoring method. As shown in Figure 17, the non-expert human achieved near-ceiling performance on Easy questions and maintained strong performance on Medium ones. Remarkably, even on Hard questions defined by consistent failure across models, the human annotator performed well above chance, achieving over 70% accuracy. This analysis underscores the significant gap between human and model capabilities on complex reasoning tasks. Even non-expert humans demonstrate strong generalization and implicit understanding, particularly in scenes that demand spatial, temporal, or motivational reasoning, highlighting the limitations of current state-of-the-art Video LMMs.
These results indicate that, while challenging, the questions are well-defined, solvable, and carefully validated requiring real reasoning capabilities.
@@ Appendix H Qualitative Results
To illustrate the complexity and diversity of implicit reasoning required in our benchmark, we present qualitative examples spanning all nine reasoning categories in Figure 22,23,24. For each example, we show the relevant video frames, question, answer choices, correct answer, and model predictions. These examples highlight a wide range of reasoning challenges -from spatial positioning and motion inference to social understanding and inherent explanation. Notably, GPT-O3 demonstrates superior performance in most cases.